NvidiaIntely (nvidiaintely):

How many oxygen atoms are in 17.63 grams of CO2? 4.823 x 1023 atoms 2.412 x 1023 atoms 4.730 x 1024 atoms 9.648 x 1023 atoms

NvidiaIntely (nvidiaintely):

@welshfella @Kevin @Will.H @brainlesslonelygirl

NvidiaIntely (nvidiaintely):

Okay, but lets see what @welshfella says.

OpenStudy (kevin):

mol = 17.63/(12+2*16) = 0.40 total atom = 6.02*10^23 x 0.40 = 2.412 * 10^23

OpenStudy (welshfella):

fraction of oxygen in CO2 = 2*atomic mass of oxygen / molecular mass of CO2 = 2*16 / 44 = 0.7273 0.7273 * 17.63 = 12.822 g O2 16 g O2 = 6.02 * 10^23 atoms 12.822 = (12.828 * 6.02 * 10^23) / 16 = 4.824 * 10^23

OpenStudy (kevin):

wait, The question is about oxygen What I calculate is molecule of CO2 1 molecule of CO2 has 2 atoms of O so, it would be 2*2.412 * 10^23 = 4.824 * 10^23

OpenStudy (will.h):

nice. i agree

OpenStudy (will.h):

with whoever is right

OpenStudy (welshfella):

hmm.. i get a different value let me check again

NvidiaIntely (nvidiaintely):

Work it out, I have all the time in the world. So don't worry :)

OpenStudy (welshfella):

Yea its A

NvidiaIntely (nvidiaintely):

Okay